OpenStudy (anonymous):

complete the square (6r^2+6r+12) = 0

OpenStudy (anonymous):

I attempted this but I checked online and got a different answer. I need some kinda of guidance to see who's right.

hartnn (hartnn):

you can first factor out 6, what u get ? and whats your attempt ?show your work till now, so that we can rectify.

OpenStudy (anonymous):

I got r=-(1/2)+or-\[\sqrt{(1/12)}\] i

hartnn (hartnn):

u just need to complete the square, right ? or u need to solve ?

OpenStudy (anonymous):

solve by completing the square

hartnn (hartnn):

so, (6r^2+6r+12) = 0 r^2 +r +2 = 0 right ?

OpenStudy (anonymous):

thats what I got

OpenStudy (anonymous):

then I moved the 2 to the other side, then took the 1 and divided by two, then squared and got r^2 + r + (1/4)= -(1/2)

hartnn (hartnn):

on right side, its -2 +1/4 = .... ? its not -1/2

OpenStudy (anonymous):

I thought you had to multiply the (1/4) by 6 since we factored that out

hartnn (hartnn):

6 is gone, forget it... x^2 + x= -2 to complete square, u juast add 1/4 on both sides.

OpenStudy (anonymous):

my teacher told me to multiply the (1/4) by the 6 that we factored out. :O

hartnn (hartnn):

i don't think that correct...

OpenStudy (anonymous):

oky, so now we have r^2 +r +(1/4)= -(7/4)

hartnn (hartnn):

yes. -7/4 is right .

hartnn (hartnn):

r^2 +r +(1/4) = (.....)^2

OpenStudy (anonymous):